# MIT-CTP/5802

### v2

@@ 3 L-GATr for Amplitude Regression
Figure 2: Left: prediction error from L-GATr and all baselines for $Z+ng$ amplitudes with increasing particle multiplicity. All networks are trained on $4\times 10^{5}$ points. Right: prediction error as a function of the training dataset size. Error bands are based on the mean and standard deviation of five random seeds. These figures are also included in Ref. [35].
Our first L-GATr case study is for amplitude regression. Partonic scattering amplitudes can be calculated exactly as a function of phase space. However, their evaluation can become very expensive if we include high-order corrections and a large number of external particles. Amplitude surrogates as part of standard event generators speed up these precision predictions [45, 46, 47, 48]. However, standard neural networks struggle to reach sufficient accuracy for realistic numbers of external particles. L-GATr uses the partial permutation symmetry of particles in the process to efficiently scale to high multiplicities, and it guarantees the Lorentz invariance of the amplitude. Extending the studies performed in Ref. [35], we demonstrate its utility for the partonic processes
$q\bar{q}\to Z+n\,g,\qquad\qquad n=1\ldots 5\;.$ (20)
@@
We generate $4\times 10^{5}$ training data points for each multiplicity up to 4 gluons with MadGraph [49]. First, we use a standard run to generate unweighted phase space points; second, we apply the standalone module to compute the squared amplitude values. To avoid divergences, we require globally
@@
  • the CGENN [14] as an equivariant graph network operating on multivectors like L-GATr.
Details about the implementation and training can be found in the Appendix A.
@@
Figure 3: Prediction error from L-GATr and selected baselines for $Z+5g$ amplitudes. Here, all networks are reduced in size and trained on $4\times 10^{4}$ points. Error bands are based on the mean and standard deviation of five random seeds.
@@
To expand on the scaling with the number of gluons, we generate $4\times 10^{4}$ points for $Z+5$ gluon productions through the pipeline described above. Given this limited number of phase space points, we reduce the size of L-GATr to $4\times 10^{4}$ parameters and the baselines to $10^{4}$ parameters, to prevent overfitting. In Fig. 3 we reproduce the L-GATr scaling behavior despite the more complex process and a less generous training.
@@ 4 L-GATr for Jet Tagging
TopoDNN [54] 0.9160 0.9720 295 $\pm$ 005
LoLa [9] 0.9290 0.9800 722 $\pm$ 017
$N$-subjettiness [55] 0.9290 0.9810 867 $\pm$ 015
@@
Table 2: Top tagging accuracy, AUC, and background rejection $1/\epsilon_{B}$ for the standard dataset [51, 2]. Lorentz-equivariant methods are indicated with an asterisk, and fine-tuning methods are separated with a horizontal line. Our error bars are based on the mean and standard deviation of five random seeds.
@@
  • PELICAN [42], an alternative equivariant graph network based on momentum invariants and permutation equivariant aggregation functions;
@@
A key ingredient for the optimization of L-GATr is the symmetry breaking prescription. For all our tests, we include two reference vectors as extra tokens: the beam direction as the $x-y$ plane bivector, $x^{B}_{12}=1$, and the time reference $x^{V}=(1,0,0,0)$, which gives the network a handle to break the symmetry down to $SO(3)$. We provide a comparison between multiple reference vector options in Tab. 3, where we test their impact on a top tagging network. From it, it is clear that both the beam direction and the time reference significantly contribute to boosting the tagging performance.
Beam Time Embedding AUC $1/\epsilon_{B}$ ($\epsilon_{S}=0.3$)
Token 0.9844 1422
$x_{3}^{V}=\pm 1$ Token 0.9850 1905
$x^{V}_{0}=1$ Token 0.9865 1923
$x_{12}^{B}=x_{13}^{B}=x_{23}^{B}=1$ $x^{V}_{0}=1$ Token 0.9863 2009
$x_{12}^{B}=1$ $x^{V}_{0}=1$ Channel 0.9865 2060
$x_{0}^{V}=1,x_{3}^{V}=\pm 1$ $x^{V}_{0}=1$ Token 0.9869 2114
$x_{3}^{V}=\pm 1$ $x^{V}_{0}=1$ Token 0.9869 2152
$x_{12}^{B}=1$ $x^{V}_{0}=1$ Token 0.9870 2240
Table 3: Symmetry breaking. We compare the L-GATr performance on the top tagging dataset from Ref. [61] using different Lorentz symmetry breaking schemes. The last line is our default used in all other tagging experiments.
@@
Table 4: Tagging accuracy, AUC, and background rejection $1/\epsilon_{B}$ for the JetClass dataset [59]. The AUC is computed as the average of all possible pairwise combinations of classes, and the acceptances are computed by comparing each signal against the background class.
@@
Table 5: Tagging accuracy, AUC, and background rejection $1/\epsilon_{B}$ on different sizes of the JetClass dataset [59]. Metrics from other models are taken from their published results [59, 60].
@@ 5 L-GATr for Event Generation
Generating LHC events is a key benchmark for neural network architectures, required for end-to-end generation, neural importance sampling, generative unfolding [64, 65, 66, 67, 68, 69], and optimal inference [70, 71]. For all these tasks we should reach per-mille-level, or at the very least percent-level accuracy on the underlying phase space density. We use the L-GATr architecture to take advantage of the approximate symmetries and to improve the scaling for increasing numbers of final state particles. Our reference process is
@@
and require two $b$-tagged jets. The events are reconstructed with a $\chi^{2}$-based algorithm [75], and identical particles are ordered by $p_{T}$. The sizes of the $t\bar{t}+n\;j$ datasets reflect the frequency of the respective processes, resulting in 9.8M, 7.2M, 3.7M, 1.5M and 480k events for $n=0\ldots 4$. We train separate networks for each multiplicity, to allow for a direct comparison between different architectures, but we emphasize that transformers can also be trained jointly on all multiplicities [76]. This is particularly useful in the case of limited training data, because transformers can transfer information across multiplicities. The results presented here were briefly discussed in Ref. [35], but without proper benchmarking.
@@
They share the generative CFM setup, which is currently the leading technique in precision generation of partonic LHC events [76] and calorimeter showers, for the latter using a vision transformer inside the CFM [80]. Our task does not require translation-equivariant representations. Therefore, we do not insert or extract features related to translational equivariance when using the E(3)-GATr, but internally it can use these representations. Details about the implementation and training can be found in the Appendix A.
Strictly speaking, the underlying problem is only symmetric under rotations around the beam axis. To not over-constrain the network we include symmetry breaking multivectors as described in Sec. 2.3. For the E(3)-GATr, we include the plane orthogonal to the beam axis as a reference multivector. For L-GATr, we also include a time reference multivector, because the distribution is not invariant under boosts along the beam axis. This step is necessary for any Lorentz-equivariant generative network, as it is not possible to construct a normalized density that is invariant under a non-compact group.
@@
Table 6: Impact of the choice of trajectory on different L-GATr networks for the CFM velocity, compared to a MLP velocity network. All networks are trained on the $t\bar{t}+0j$ dataset.
@@
In practice, we encounter large jacobians from the logarithm transformations in the above relation for $v_{x_{m}},v_{x_{p}}$ due to small values of the jet mass $m$ and the relative transverse momentum $p_{T}-p_{T,\mathrm{min}}$, leading to unstable training. To avoid this obstacle, we add a fourth step to the procedure, where we overwrite the two problematic velocity components with scalar outputs of the L-GATr network. This adds an additional redundant source of symmetry breaking to the reference multivectors discussed above.
@@
Figure 6: Marginal distributions for $t\bar{t}+1,2,3,4$ jets (top to bottom). We do not show E(3)-GATr results, as they are very similar to the standard transformer. The three panels in the bottom row are also included in Ref. [35].
Figure 7: Performance of generative networks in terms of a negative log-likelihood over the events (left) and a trained classifier AUC (right). In the top row we show the scaling with the size of the training dataset for the mixed $t\bar{t}+n$ jet dataset, in the bottom row the scaling with the number of particles in the final state. The MLP, Transformer and L-GATr metrics were already discussed in Ref. [35].
We compare a set of 1-dimensional distributions from the different generators in Fig. 6. Observables like $p_{T,b}$, that are part of the phase space parametrization Eq. (32), are easily learned by all networks because of our choice of target trajectories in Eq. (29). All other observables appear s correlations and are harder to learn. L-GATr outperforms the baselines across all distributions. Especially angular correlations benefit from the equivariance encoded in the L-GATr architecture, enabling percent-level precision in these variables for the first time. The main weakness of all architectures are the intermediate top mass poles, requiring the correlation of three external 4-vectors.
To analyze scaling properties, we use the negative log-likelihood evaluated on the events and the AUC of a neural classifier. In Fig. 7 we find a clear performance increase with increasing symmetry awareness, from the unstructured MLP over the permutation-equivariant transformer to the rotation-equivariant GATr and the Lorentz-equivariant L-GATr. In particular, the superior L-GATr performance mainly originates from boost-equivariance, as the rotation-equivariant E(3)-GATr performs only marginally better than the plain transformer. This might come as a surprise, as we allow L-GATr to break this boost equivariance using reference multivectors. This implies that enforcing equivariance in the architecture and then allowing the network to break it with reference multivectors outperforms standard non-equivariant networks.
@@ 6 Outlook
Modern ML at the LHC has developed from mostly concept development to the first applications in experiment and theory. For these applications, performance is the main goal, leading us to the question how we can train the most precise neural networks on a large, but nevertheless limited training dataset. In LHC physics, we are in the lucky situation that we can use the known structure of the phase space. It rests on a complex system of symmetries, the leading one being the Lorentz symmetry.
@@
L-GATr is a versatile equivariant transformer that constructs such a Lorentz representation for regression, classification, and generation networks. For amplitude regression, a key step in speeding up loop amplitudes in event generators, L-GATr shows the best performance for more than three particles in the final state, thanks to its superior data efficiency, leading to an improved scaling with the phase space dimensionality. For subjet tagging, L-GATr combines the benefit of equivariance with pre-training on large datasets and is at least on par with the best available subjet tagger. Finally, the combination of L-GATr with CFM generator faithfully reproduces the phase space distribution of top pair production with up to four jets better than all other CFM setups. We look forward to further applications of L-GATr at the LHC, as well as generalizations to incorporate additional domain knowledge from collider physics.
@@
L-GATr is available at https://github.com/heidelberg-hepml/lorentz-gatr as part of the public Heidelberg hep-ml code and tutorial library.

### v3

@@ 3 L-GATr for Amplitude Regression
Figure 2: Left: prediction error from L-GATr and all baselines for $Z+ng$ amplitudes with increasing particle multiplicity. All networks are trained on $4\times 10^{5}$ points. Right: prediction error as a function of the training dataset size. Error bands are based on the mean and standard deviation of five random seeds affecting network weight initialization. These figures are also included in Ref. [35].
Our first L-GATr case study is for amplitude regression. Partonic scattering amplitudes can be calculated exactly as a function of phase space. However, their evaluation can become very expensive if we include high-order corrections and a large number of external particles. Amplitude surrogates as part of standard event generators speed up these precision predictions [45, 46, 47, 48]. However, standard neural networks struggle to reach sufficient accuracy for a high amount of external particles. L-GATr uses the partial permutation symmetry of particles in the processes to efficiently scale to high multiplicities, and it guarantees the exact Lorentz invariance of the amplitude. Extending the studies performed in Ref. [35], we demonstrate its utility for the partonic processes
$q\bar{q}\to Z+n\,g,\qquad\qquad n=1\ldots 4\;.$ (20)
@@
We generate $4\times 10^{5}$ training data points for each multiplicity up to 4 gluons with MadGraph [49] at LO in QCD. First, we use a standard run to generate unweighted phase space points; second, we apply the standalone module to compute the squared amplitude values. To avoid divergences, we require globally
@@
  • CGENN [14] as an equivariant graph network operating on multivectors like L-GATr.
All of these networks are trained with 4-momenta inputs except DSI, which also takes pairwise Minkowski products of 4-momenta as inputs. Further details about the implementation and training can be found in the Appendix A.
@@
Figure 3: Prediction error from L-GATr and selected baselines for $Z+5g$ amplitudes. Here, all networks are reduced in size and trained on $4\times 10^{4}$ points. Error bands are based on the mean and standard deviation of five random seeds affecting network weight initialization.
@@
To expand on the scaling with the number of gluons, we generate $4\times 10^{4}$ points for $Z+5$ gluon productions through the pipeline described above. Given this limited number of phase space points, we reduce the size of L-GATr to $4\times 10^{4}$ parameters and the baselines to $10^{4}$ parameters, to prevent overfitting. For the sake of comparison with the smaller multiplicities, we also reduce the size of the rest of the datasets to $4\times 10^{4}$ points. In Fig. 3 we reproduce the L-GATr scaling behavior despite the more complex process and a less generous training. Due to optimization challenges caused by the small data regime, we only compare L-GATr to the transformer and DSI, representing an attention-based network and a symmetry-enforcing network respectively.
@@ 4 L-GATr for Jet Tagging
TopoDNN [54, 2] 0.9160 0.9720 295 $\pm$ 005
LoLa [9, 2] 0.9290 0.9800 722 $\pm$ 017
NN-subjettiness [55, 2] 0.9290 0.9810 867 $\pm$ 015
@@
Table 2: Top tagging accuracy, AUC, and background rejection $1/\epsilon_{B}$ at different signal efficiencies $\epsilon_{S}$ for the standard dataset [51, 2]. Lorentz-equivariant methods are indicated with an asterisk, and fine-tuning methods are separated with a horizontal line. Entries with two citations correspond to architectures whose quoted results were presented in the global analysis from Ref. [2]. Our error bars are based on the mean and standard deviation of five random seeds affecting network weight initialization.
@@
  • PELICAN [42], an equivariant graph network based on momentum invariants and permutation equivariant aggregation functions;
@@
A key ingredient for the optimization of L-GATr is the symmetry breaking prescription. For all our tests, we include reference vectors as extra tokens or channels: the beam direction as the xyx-y plane bivector, $x^{B}_{12}=1$, and the time reference $x^{V}=(1,0,0,0)$, which gives the network a handle to break the symmetry down to $SO(3)$. We provide a comparison between multiple reference vector options in Tab. 3, where we test their impact on a top tagging network. In it, we explore four aspects of our symmetry breaking approach: the time reference addition, the beam embedding format, the addition of extra kinematic features, and the inclusion of the reference vectors as either extra tokens or extra channels for each token. We conclude from these results that including both the beam direction and the time reference significantly contributes to boosting the tagging performance. The best performing setup according to the area under the curve (AUC) is used for all tagging experiments. The background rejection is also considered for this choice, but we do not use it as a deciding factor due to its larger uncertainty.
Beam Time Embedding Extra features AUC $1/\epsilon_{B}$ ($\epsilon_{S}=0.3$)
Token 0.9846 $\pm$ 0.0002 1494 $\pm$ 62
$x_{3}^{V}=\pm 1$ Token 0.9854 $\pm$ 0.0005 1684 $\pm$ 221
$x_{12}^{B}=x_{13}^{B}=x_{23}^{B}=1$ $x^{V}_{0}=1$ Token 0.9864 $\pm$ 0.0002 2076 $\pm$ 75
$x^{V}_{0}=1$ Token 0.9865 $\pm$ 0.0001 2179 $\pm$ 123
$x_{12}^{B}=1$ $x^{V}_{0}=1$ Channel 0.9866 $\pm$ 0.0001 2150 $\pm$ 82
$x_{0}^{V}=1,x_{3}^{V}=\pm 1$ $x^{V}_{0}=1$ Token 0.9869 $\pm$ 0.0001 2282 $\pm$ 174
$x_{3}^{V}=\pm 1$ $x^{V}_{0}=1$ Token 0.9869 $\pm$ 0.0001 2293 $\pm$ 141
Token 0.9869 $\pm$ 0.0001 2179 $\pm$ 156
$x_{12}^{B}=1$ $x^{V}_{0}=1$ Token 0.9870 $\pm$ 0.0001 2173 $\pm$ 83
$x_{12}^{B}=1$ $x^{V}_{0}=1$ Token 0.9870 $\pm$ 0.0001 2240 $\pm$ 70
Table 3: Symmetry breaking for tagging. We compare the L-GATr performance on the top tagging dataset from Ref. [61] using different Lorentz symmetry breaking schemes. Our error bars are calculated as the standard deviation of three random seeds affecting network weight initialization. The last line is our default used in all other tagging experiments.
@@
Table 4: Tagging accuracy, AUC, and background rejections $1/\epsilon_{B}$ for the JetClass dataset [59]. The AUC is computed as the average of all possible pairwise combinations of classes, and the acceptances are computed by comparing each signal against the background class.
@@
Table 5: Tagging accuracy, AUC, and background rejections $1/\epsilon_{B}$ on different sizes of the JetClass dataset [59]. Metrics from other models are taken from their published results [59, 60].
@@ 5 L-GATr for Event Generation
Generating LHC events is a key benchmark for neural network architectures, required for end-to-end generation, neural importance sampling, generative unfolding [64, 65, 66, 67, 68, 69], and optimal inference [70, 71]. For all these tasks we should reach per-mille-level (or at the very least percent-level) accuracy on the underlying phase space density. We use the L-GATr architecture to take advantage of the approximate symmetries and to improve the scaling for increasing numbers of final state particles. Our reference process is
@@
and require two $b$-tagged jets. The events are reconstructed with a $\chi^{2}$-based algorithm [75], and identical particles are ordered by $p_{T}$. The sizes of the $t\bar{t}+n\;j$ datasets reflect the frequency of the respective processes, resulting in 9.8M, 7.2M, 3.7M, 1.5M and 480k events for $n=0\ldots 4$. The results presented here were briefly discussed in Ref. [35], but without proper benchmarking.
@@
They share the generative CFM setup, which is currently the leading technique in precision generation of partonic LHC events [77] and calorimeter showers, with the latter using a vision transformer as the CFM neural network [80]. Our task does not require translation-equivariant representations. Therefore, we do not insert or extract features related to translational equivariance when using the E(3)-GATr, even though it can use these representations internally. Details about the implementation and training can be found in Appendix A.
Strictly speaking, the underlying process is only symmetric under rotations around the beam axis. To not over-constrain the network, we include symmetry breaking multivectors as described in Sec. 2.3. For the E(3)-GATr, we include the plane orthogonal to the beam axis as a reference multivector. For L-GATr, we also include a time reference multivector, because the distribution is not invariant under boosts along the beam axis. This step is necessary for any Lorentz-equivariant generative network, as it is not possible to construct a normalized density that is invariant under a non-compact group.
@@
Table 6: Impact of the choice of trajectory on different L-GATr networks for the CFM velocity, compared to a MLP velocity network. All networks are trained on the $t\bar{t}+0j$ dataset. Error bars are based on the mean and standard deviation of three random seeds and include effects from network weight initialization, classifier trainings and likelihood evaluation based on an differential equation solver.
@@
In practice, we encounter large Jacobian matrix components from the logarithm transformations in the above relation for $v_{x_{m}},v_{x_{p}}$ due to small values of the jet mass $m$ and the relative transverse momentum $p_{T}-p_{T,\mathrm{min}}$, leading to unstable training. To avoid this obstacle, we add a fourth step to the procedure, where we overwrite the two problematic velocity components with scalar outputs of the L-GATr network. This adds an additional redundant source of symmetry breaking to the reference multivectors discussed above.
@@
In Tab. 7, we compare the impact of various symmetry breaking schemes on event generation. As in Tab. 3, reference multivectors are incorporated as extra channels or tokens. Once again, we observe that including reference multivectors for both beam and time directions is essential; omitting them leads to inferior performance of the L-GATr generator compared to a non-equivariant transformer. However, the specific representation of the beam reference multivector is less critical here than in the top-tagging case, as all tested representations yield consistent results within statistical uncertainty. Including reference multivectors as extra tokens in this case adds significant computational cost due to the comparably low number of particles, so we choose to include them as extra channels in our default setting.
Beam Time Embedding NLL AUC
Channel 29.36 $\pm$ 1.34 0.996 $\pm$ 0.001
$x^{V}_{0}=1$ Channel -18.77 $\pm$ 0.78 0.990 $\pm$ 0.001
$x_{12}^{B}=1$ Channel -32.04 $\pm$ 0.03 0.711 $\pm$ 0.001
$x_{12}^{B}=1$ $x^{V}_{0}=1$ Token -32.60 $\pm$ 0.03 0.523 $\pm$ 0.014
$x_{0}^{V}=\sqrt{2},x_{3}^{V}=\pm 1$ $x^{V}_{0}=1$ Channel -32.66 $\pm$ 0.02 0.510 $\pm$ 0.001
$x_{3}^{V}=\pm 1$ $x^{V}_{0}=1$ Channel -32.65 $\pm$ 0.01 0.510 $\pm$ 0.001
$x_{0}^{V}=1,x_{3}^{V}=\pm 1$ $x^{V}_{0}=1$ Channel -32.64 $\pm$ 0.04 0.513 $\pm$ 0.005
$x_{12}^{B}=1$ $x^{V}_{0}=1$ Channel -32.64 $\pm$ 0.02 0.514 $\pm$ 0.006
Table 7: Symmetry breaking for event generation. We compare the L-GATr performance on the $t\bar{t}+0j$ dataset using different Lorentz symmetry breaking schemes. The last line is our default used in all other generation experiments. Error bars are based on the mean and standard deviation of three random seeds and include effects from network weight initialization, classifier trainings and likelihood evaluation based on an differential equation solver.
@@
Figure 6: Overview of marginal distributions for $t\bar{t}+1,2,3,4$ jets (top to bottom). We do not show E(3)-GATr results, as they are very similar to the standard transformer. The three panels in the bottom row are also included in Ref. [35].
Figure 7: Performance of generative networks in terms of a negative log-likelihood over the events (left) and a trained classifier AUC (right). In the top row we show the scaling with the size of the training dataset for the mixed $t\bar{t}+n$ jet dataset, in the bottom row the scaling with the number of particles in the final state. The MLP, Transformer and L-GATr metrics were already discussed in Ref. [35]. Error bands are based on the mean and standard deviation of five random seeds and include effects from network weight initialization, classifier trainings and likelihood evaluation based on an differential equation solver.
We present a limited set of 1-dimensional distributions from the different generators in Fig. 6. Observables like $p_{T,b}$, which are part of the phase space parametrization Eq. (32), are easily learned by all networks because of our choice of target trajectories in Eq. (29). All other observables appear as correlations and are harder to learn. L-GATr outperforms the baselines across all distributions. Angular correlations especially benefit from the equivariance encoded in the L-GATr architecture, enabling percent-level precision in these variables for the first time. The main weakness of all architectures are the intermediate top mass poles, requiring the correlation of three external 4-vectors.
To analyze scaling properties, we use the negative log-likelihood evaluated on the events and the AUC of a neural classifier. In Fig. 7 we find a clear performance improvement as symmetry awareness increases, from the unstructured MLP over the permutation-equivariant transformer to the rotation-equivariant GATr and the Lorentz-equivariant L-GATr. In particular, the superior L-GATr performance mainly originates from boost-equivariance, as the rotation-equivariant E(3)-GATr performs only marginally better than the plain transformer. This might come as a surprise, as we allow L-GATr to break this boost equivariance using reference multivectors. This implies that enforcing equivariance in the architecture and then allowing the network to break it with reference multivectors outperforms standard non-equivariant networks.
@@ 6 Outlook
Modern ML at the LHC has evolved from mostly concept development to the first applications in experiment and theory. For these applications, performance is the main goal, leading us to the question of how we can train the most precise neural networks on a large, but nevertheless limited training dataset. In LHC physics, we are in the lucky situation that we can use the known structure of the phase space. It rests on a complex system of symmetries, the leading one being the Lorentz symmetry.
@@
L-GATr is a versatile equivariant transformer that constructs such a Lorentz representation for regression, classification, and generation networks. For amplitude regression, a key step in speeding up loop amplitudes in event generators, L-GATr shows the best performance for more than three particles in the final state, thanks to its superior data efficiency, leading to an improved scaling with the phase space dimensionality. For subjet tagging, L-GATr combines the benefit of equivariance with pre-training on large datasets and is at least on par with the best available subjet tagger. Finally, the combination of L-GATr with a CFM generator faithfully reproduces the phase space distribution of top pair production with up to four jets better than all other CFM setups. We look forward to further applications of L-GATr at the LHC, as well as generalizations to incorporate additional domain knowledge from collider physics.
@@
L-GATr is available at https://github.com/heidelberg-hepml/lgatr as part of the public Heidelberg hep-ml code and tutorial library. The results in this publication can be reproduced with https://github.com/heidelberg-hepml/lorentz-gatr.